The sterilization tank contains 100 gallons of a solution of 2% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 5% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?



Answer by Theo(13342)

we have 100 gallons of 2% bleach

we need 100 gallon of 5% bleach.

100 * .02 = 2 gallons of bleach in the original mixture.

we need 100 * .05 = 5 gallons of bleach in the new mixture.

let x represent the number of gallons that have to be replaced.

we want to have 5 gallons in the new mixture.

we have 2 gallons in the original mixture.

we will need to add x gallons of pure bleach.

in the process we will be removing .02*x gallons of bleach from the original mixture.

the formula becomes:

5 = 2 - .02*x + x

where:

5 is the number of gallons of bleach we want in the new mixture.
2 is the number of gallons of bleach in the original mixture.
x is the number of gallons of pure bleach we will be adding.
.02*x is the number of gallons of pure bleach we will be removing.

the formula simplifies to:

5 = 2 + .98x

subtract 2 from both sides of this equation to get:

3 = .98x

divide both sides of this equation by .98 to get:

x = 3/.98 = 3.06122449

they need to replace 3.06122449 gallons of a 2% mixture with 3.06122449 gallons of pure bleach to reach the required concentration.

assume they remove 3.06122449 gallons of the 2% mixture.

they now have 96.93877551 gallons of 2% mixture.

this means they have .02 * 96.93877551 = 1.93877551 gallons of bleach.

they add 3.06122449 gallons of pure bleach.

their total gallons becomes 96.93877551 + 3.06122449 = 100 gallons.

their bleach content becomes 1.93877551 + 3.06122449 = 5 gallons.

they now have a 5% solution.

they needed to replace 3.06122449 gallons of the 2% mixture with 3.06122449 gallons of pure bleach.